Saddle Inc. has two types of handbags: standard and custom. The controller has decided to use a plantwide overhead rate based on direct labor costs. The president has heard of activity-based costing and wants to see how the results would differ if this system were used. Two activity cost pools were developed: machining and machine setup. Presented below is information related to the company’s operations.

Standard

Custom

Direct labor costs $40,000 $110,000
Machine hours 1,450 1,460
Setup hours 103 370


Total estimated overhead costs are $295,000. Overhead cost allocated to the machining activity cost pool is $196,000, and $99,000 is allocated to the machine setup activity cost pool.

Compute the overhead rate using the traditional (plantwide) approach. (Round answer to 2 decimal places, e.g. 12.25.)

Predetermined overhead rate

enter the predetermined overhead rate in percentages of direct labor cost rounded to 2 decimal places

% of direct labor cost

eTextbook and Media

  

  

Compute the overhead rates using the activity-based costing approach. (Round answers to 2 decimal places, e.g. 12.25.)

Machining

$enter a dollar amount per machine hour rounded to 2 decimal places

per machine hour
Machine setup

$enter a dollar amount per setup hour rounded to 2 decimal places

per setup hour

eTextbook and Media

  

  

Determine the difference in allocation between the two approaches. (Round answers to 0 decimal places, e.g. 1,225.)

Traditional costing
Standard

$enter a dollar amount rounded to 0 decimal places

Custom

$enter a dollar amount rounded to 0 decimal places

Activity-based costing
Standard

$enter a dollar amount rounded to 0 decimal places

Custom

$enter a dollar amount rounded to 0 decimal places
Answer #1
1
Total estimated overhead 295000
Divide by Total Direct labor costs 150000 =40000+110000
Predetermined overhead rate 196.67 % of direct labor cost
2
Machining 67.35 per machine hour =196000/(1450+1460)
Machine setup 209.30 per setup hour =99000/(103+370)
3
Traditional costing:
Standard 78668 =40000*196.67%
Custom 216337 =110000*196.67%
Activity-based costing :
Standard 119215 or 119216 =(1450*67.35)+(103*209.30)
Custom 175772 =(1460*67.35)+(370*209.30)